OpenStudy (anonymous):

Factor Each Trinomial: a^2 - 5^2 - 20 a^2 +12ab + 27b^2

OpenStudy (anonymous):

Help plz!!

OpenStudy (anonymous):

a^2 + 12ab + 27b^2 = (a + 9b)(a + 3b)

OpenStudy (anonymous):

Merci

OpenStudy (anonymous):

@myko NO. \[\LARGE (a+4)(a-5)=a^2-5a+4a-20=a^2-a-20\] not.. \[\LARGE a^2-5^2-20\] because that's what she's asking!

OpenStudy (anonymous):

Oo. Sry. Didn't read question carefully @Kreshnik. (a-sqrt45)(a+sqrt45)

OpenStudy (anonymous):

Correct. Well done ! xD